Would you please tell me the steps to solve this equation?
+-45q+to+the+11th+power+divided+by+5q+to+the+7th+power+
Thank you.
Allison

Answer by checkley79(3341)

-45q^11/5q^7
-45/5=-9
q^11/q^7=q^4
-9q^4 ans.